# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

### v1

@@ Abstract
Political actors frequently manipulate redistricting plans to gain electoral advantages, a process commonly known as gerrymandering. To address this problem, several states have implemented institutional reforms including the establishment of map-drawing commissions. It is difficult to assess the impact of such reforms because each state structures bundles of complex rules in different ways. We propose to model redistricting processes as a sequential game. The equilibrium solution to the game summarizes multi-step institutional interactions as a single dimensional score. This score measures the leeway political actors have over the partisan lean of the final plan. Using a differences-in-differences design, we demonstrate that reforms reduce partisan bias and increase competitiveness when they constrain partisan actors. We perform a counterfactual policy analysis to estimate the partisan effects of enacting recent institutional reforms nationwide. We find that instituting redistricting commissions generally reduces the current Republican advantage, but Michigan-style reform would yield a much greater pro-Democratic effect than types of redistricting commissions adopted in Ohio and New York.
@@ 1 Introduction
Democratic institutions play a crucial role in preventing political actors from pursuing policies that prioritize their own interests over the broader public good (Federalist papers 53 and 54: Madison, 1788a, ; Madison, 1788b, ). As political polarization increases and democratic norms erode, scholars have recognized the growing importance of institutional constraints on policy makers (Levitsky and Ziblatt,, 2023; Little and Meng,, 2023; McCarty,, 2019). To safeguard institutions, advocates have suggested reforms which insulate democratic processes from partisan control.222For example, see People Not Politicians Oregon or Ohio’s Citizens Not Politicians. For example, the Electoral Count Reform and Presidential Transition Improvement Act of 2022 represents an effort to make it more difficult for partisan actors to manipulate the presidential electoral certification process.
@@
This paper studies American congressional redistricting, a political process often exploited by partisan actors to enact districting plans that favor their own party. This manipulation, known as partisan gerrymandering, has been widespread in the past two redistricting cycles (Kenny et al.,, 2024, 2023; Warshaw et al.,, 2022). Redistricting plans that disproportionately favor a certain party can limit how responsive a party’s share of seats in the legislature are to changes in its vote share, and can reduce the electoral power of racial minorities (e.g. Canon,, 2022; Canon and Race,, 1999; Grofman and Handley,, 1991; Polsby and Popper,, 1991).
@@
Reform efforts to limit gerrymandering are often designed to constrain partisan map drawers. They include the establishment of independent map-drawing commissions and the introduction of court oversight over proposed plans (Cain,, 2012). Estimating the causal impact of these institutional reforms, however, is challenging for three reasons. First, as in any observational study, states that adopt redistricting reforms may differ in important ways from states that do not. Second, the reforms themselves intervene on complex and interrelated institutional processes, which differ widely even among the reforming states. Third, the outcome of the redistricting process is a districting plan, whose partisan features may be confounded by other factors such as a state’s geography and demographics that are possibly correlated with reform efforts. Thus, both the treatment and outcome variables of our study are multifaceted and multidimensional. This means that most existing causal methodology, which considers a binary treatment and univariate outcome, is not readily applicable.
@@
We estimate the causal effects of redistricting reforms by combining a formal theoretic approach with statistical methods of causal inference. To reduce the differing districting procedures across states to a single theoretically informed parameter, we model redistricting reform as a zero-sum sequential game. We collect the information about each state’s relevant laws, starting from the initial map drawer (i.e., the redistricting commission or legislature) and following through varying stalemate processes and opportunities for court intervention. This new dataset is used to structure the players and available moves in the game. Analyzing the game allows us to estimate the ability of partisan players to maximize the partisan lean of a redistricting plan. The Nash equilibrium of the game is a measure of the “leeway” that a single party has over the final redistricting plan.
@@
We then use these leeway measures as a continuous treatment variable in a differences-in-differences (DiD) design applied to the 2010 and 2020 redistricting cycles. This addresses potential confounding by comparing changes in states that have enacted reforms to those in similar states that have not, relying on a parallel trends assumption (Callaway et al.,, 2024). We estimate how changes in map drawer’s leeway may influence resulting plans in terms of traditional measures of partisan fairness, such as the efficiency gap (Stephanopoulos and McGhee,, 2015) as well as nonpartisan “good-governance” outcomes, like the responsiveness to swings in partisan preferences.
@@
Lastly, we separate the causal effect of state-specific institutional reforms from the change in political geography over time. To do this, we generate a sample of alternative redistricting plans via a simulation algorithm (McCartan and Imai,, 2023) following each state’s rules and political geography, but without regard to any partisan information (McCartan et al.,, 2022; Kenny et al.,, 2024). We use these simulated plans as a nonpartisan baseline for both 2010 and 2020 redistricting cycles, and compute the difference in outcome variables between the enacted and simulated plans. This adjusts for state-specific changes in political geography, making the parallel trends assumption more plausible.
@@
We find that more restrictive redistricting processes reduce partisan bias by constraining map drawers. For example, changing from a single-party legislature to an independent commission leads to about an reduction in excess seats of about 0.8 seats. We find smaller but positive effects of constraining reforms on electoral responsiveness. We estimate that a similar change from legislature to commission would increase the share of competitive seats in an average state from 25% to 32%.
@@
One advantage of our methodological approach is the ability to conduct a counterfactual analysis of institutional reforms. We investigate how enacting recent procedural reforms nationwide could reduce widespread partisan gerrymandering. We quantify how the partisan bias and responsiveness of adopted plans could counterfactually change if all states nationwide adopted three kinds of commission structures currently enacted in several states: (1) an Ohio-style approach that requires supermajorities and uses a bipartisan backup commission, (2) a New York-style commission with a nonpartisan map drawer but several partisan veto points later in the process, and (3) a Michigan-style commission with a nonpartisan commission drawer, no partisan veto points, and the potential for court review.
@@
We find that commissions can generally reduce the existing Republican bias but the details of commission structure matter. In particular, unlike reforms adopted in Ohio and New York, a Michigan-style non-partisan commission has no partisan veto points. We find that implementing Michigan-style reforms nationwide leads to an additional 22 Democratic seats, on average. Michigan-style reforms also increase electoral responsiveness whereas those adopted in New York and Ohio do not.
@@
Beyond the substantive findings, this paper advances a new methodology for studying redistricting reform, one which we hope will translate to the study of other institutional systems as well. Redistricting reform efforts have produced diverse institutional changes across states. While most scholars have simply classified reforms into different categories (Cain,, 2012; Edwards et al.,, 2017; Warshaw et al.,, 2022; Nelson,, 2023), such an approach may miss important differences in these institutional changes.
@@
In contrast, we use formal modeling to place these complex institutional characteristics on a continuous univariate scale to summarize how they constrain partisan actors. This theoretically driven approach, which we demonstrate predicts the empirical patterns well, makes it possible to apply the difference-in-differences strategy to our complex setting for credible causal inference.
@@
While our work estimates the causal effects of redistricting reform as a whole, much of existing causal analysis has focused on one part of the process, such as registration patterns after the Shelby decision (Komisarchik and White,, 2021) and turnout in packed districts (Fraga et al.,, 2022). Other studies have reported positive associations between the adoption of commissions and competitiveness (Carson and Crespin,, 2004; Carson et al.,, 2014; Best et al.,, 2021; Nelson,, 2023) as well as a negative association between commissions and incumbent protection (Nelson,, 2023). These studies, however, do not adjust for various confounding factors considered in our analysis.
@@
Beyond redistricting, our methodological approach can be seen as a general strategy for causal analysis of complex institutional reforms. By combining a game-theoretic treatment model with causal inference methods, we are able to leverage the strengths of these two approaches. A game-theoretic approach is a substantively effective way to model formal institutions with specific rules, making it possible to map multidimensional policies to a univariate summary in a theoretically-informed way. Once this summarization of various institutional reforms is done, we can apply standard causal inference methods to estimate the effects of counterfactual policies by mapping them directly to this univariate treatment variable.
@@
Our work relates to a broader methodological literature in the social sciences. In economics, for example, Chetty, (2009) advocates the combined use of structural and reduced-form approaches via a sufficient statistic in the manner similar to our analysis. In sociology, Lundberg et al., (2021) emphasize the importance of theoretically informed quantities of interest in causal analysis. Finally, in political science, Canen and Ramsay, (2023) call for the integration of rigorous theoretical and empirical approaches in causal research. We demonstrate how such an analysis can be done in the estimation of causal effects of redistricting reforms.
@@ 2 Overview of Redistricting Processes and Reforms
Before introducing our methodological framework, we provide a brief overview of redistricting processes and reforms. Every decade following the U.S. Census, states and localities use many different procedures to redraw their legislative district boundaries. For example, states differ on whether legislatures or independent redistricting commissions propose initial plans for new Congressional district maps. If these actors fail to produce a plan, state laws further vary on how these stalemates are handled. Some states pass responsibilities to a court (e.g., Virginia), a backup commission (e.g., Ohio), or a group of state party leaders (e.g., Iowa). Even plans that pass the proposal step can face a veto from other actors like the governor or state legislature.
As explained in Section 3.1 below, we collect information about the redistricting process used in each state for the 2010 and 2020 cycles. Figure 1 summarizes this dataset and illustrates the diversity of 2020 redistricting procedures across states. For example, 29 states drew initial plans in their legislature, while 15 states used an independent redistricting commission instead. There is significant variation in procedures following this initial draw, with seven distinct veto mechanisms across all the states. Much of this variation originates from state-specific policy campaigns that have adopted widely different goals driven by local actors (Keena et al.,, 2021). For example, California created a redistricting commission that has the power of drawing congressional district boundaries through a series of ballot propositions in the early 2000s. In recent years, many states have also implemented redistricting reforms often driven by desires to limit partisan influence. For example, Michigan voters approved a constitutional amendment to establish an independent redistricting commission in 2018.
@@
While these simplifications make the study of complex procedures tractable, they risk oversimplifying complex procedures in two ways. First, attributing institutional outcomes to a single actor ignores the fact that the final redistricting plan is produced through a series of steps. Take, for example, the New York redistricting process in 2020. An independent bipartisan commission had the power to propose an initial plan, but failed to agree on a single plan. State law required the stalemated process to move to the legislature, which adopted a plan that Republican and civil rights groups criticized as a Democratic gerrymander. These groups challenged the adopted plan in a series of lawsuits, and in 2022 the New York Court of Appeals struck down the plan and tasked a court-appointed special master with drawing a remedial plan. Though the final plan scores well on quantitative fairness metrics (Kenny et al.,, 2023), simply classifying a “court” as the sole map drawer for New York in 2020 overlooks the partisan interests involved, failing to capture the complexity of the procedures that led to the final plan.
Second, simple classification schemes can overlook strategic interactions, where the behavior of certain actors depend on the presence or characteristics of others. For example, a commission might draw a different map if it knew the map could later be reviewed by a court. Or, the potential of a governor’s veto may limit the likelihood of partisan gerrymander by the state legislature, but not in cases where the governor and legislative majority share a partisan affiliation. More detailed coding of procedural schemes can account for these interaction effects and increase realism, but will necessarily decrease statistical power, making it more difficult to estimate causal effects.
Thus, we face a methodological dilemma: while some simplification of institutional features is necessary, common approaches to doing so obscure critical characteristics of redistricting processes. In the next section, we propose a theoretically-grounded approach that models redistricting as a zero-sum sequential game and uses its Nash equilibrium as a treatment variable summarizing complex institutions.
@@ 3 A Theoretical Model of Institutional Leeway
To study the impact of redistricting reforms and processes, we first develop a standardized set of 14 institutional features that capture the most important actors in the congressional redistricting processes of the states. We collect the data on these features so that it is possible to compare redistricting processes consistently across states, despite the differences in laws and bodies governing redistricting in each state.
Next, we develop a sequential redistricting game to summarize these features in a theoretically informed manner. Specifically, we use the Nash equilibrium of this game as a one-dimensional measure of the institutional leeway political actors have over the partisan lean of the resulting redistricting plan. Finally, we empirically validate the proposed measure by demonstrating that it is not particularly sensitive to model specification and predicts redistricting outcomes well.

### v2

@@ Abstract
Political actors often manipulate redistricting plans to gain electoral advantages, a process known as gerrymandering. Several states have implemented institutional reforms to address this problem, such as establishing map-drawing commissions. Estimating the impact of such reforms is challenging because each state structures bundles of rules in different ways. We model redistricting as a sequential game, where each state’s equilibrium solution summarizes multi-step institutional interactions as a single-dimensional score. We argue this score measures the leeway political actors have over the partisan lean of the final plan. Using a differences-in-differences design, we demonstrate that reforms reduce partisan bias and increase competitiveness when they constrain partisan actors. We perform a counterfactual policy analysis to estimate the effects of enacting recent reforms nationwide. Though commissions generally reduce bias, reforms that restrict partisan actors in multiple ways like removing veto points (Michigan) are much more effective than commissions where parties retain some control (Ohio).
@@ 1 Introduction
Democratic institutions play a crucial role in preventing political actors from pursuing policies that prioritize their own interests over the broader public good (Federalist papers 53 and 54: Madison, 1788a, ; Madison, 1788b, ). Further, increasing political polarization and the erosion of democratic norms in many countries have encouraged more recent attention to placing additional institutional constraints on policy makers (Levitsky and Ziblatt,, 2023; Little and Meng,, 2023; McCarty,, 2019). To safeguard institutions, advocates have suggested reforms that insulate democratic processes from partisan control.222For example, see People Not Politicians Oregon or Ohio’s Citizens Not Politicians. For example, the Electoral Count Reform and Presidential Transition Improvement Act of 2022 represents an effort to make it more difficult for partisan actors to manipulate the presidential electoral certification process in the United States.
@@
We study reform in American congressional redistricting, a political process often exploited by partisan actors to enact districting plans that favor their own party. This manipulation, known as partisan gerrymandering, has been widespread in the past two redistricting cycles (Kenny et al.,, 2024, 2023; Warshaw et al.,, 2022). Redistricting plans that disproportionately favor a certain party can limit how responsive a party’s share of seats in the legislature is to changes in its vote share and can reduce the electoral power of racial minorities (e.g. Canon,, 2022; Canon and Race,, 1999; Grofman and Handley,, 1991; Polsby and Popper,, 1991).
@@
Reform efforts to limit gerrymandering are often designed to constrain partisan map drawers. They include the establishment of independent map-drawing commissions and the introduction of court oversight over proposed plans (Cain,, 2012). Estimating the causal impact of these institutional reforms, however, is challenging for three reasons. The first is the problem of treatment complexity. Redistricting reform efforts must intervene in a complex and multidimensional process. The specific rules governing each state’s redistricting process differ in many ways, including who proposes initial maps and whether or not courts can intervene. Second, states that adopt redistricting reforms may differ in both observable and unobservable ways from states that do not. This selection bias problem is common to any observational study and must be addressed to estimate credible causal effects.
Finally, the outcome of the institutional process — a redistricting plan — is also complex. For example, partisan features may be confounded by other factors such as a state’s geography and demographics (e.g., Cottrell,, 2019). If a large number of Democratic voters live in cities, any redistricting plan with compact districts may end up creating a small number of heavily Democratic-leaning districts rather than efficiently allocating Democratic votes across more districts. We must overcome this outcome complexity to accurately measure the partisan bias of each enacted plan.
@@
In this paper, we propose a new methodological approach that addresses the above challenges to study the causal impact of redistricting reform. To deal with the treatment complexity, we first collect and standardize information about each state’s relevant laws, starting from the initial map drawer (i.e., the redistricting commission or legislature) and following through varying stalemate processes and opportunities for court intervention. We then reduce the differing districting procedures across states to a single theoretically informed parameter by modeling redistricting reform as a zero-sum sequential game. We use our original dataset of institutional procedures to characterize the players and available moves in the game. Analyzing the game allows us to measure the ability of partisan players to maximize the partisan lean of a redistricting plan. The Nash equilibrium of the game is a measure of the “leeway” that a single party has over the final redistricting plan.
We use two variations of each state’s game to produce two measures of leeway. The first is the realized leeway, which uses the observed parties of the players to compute the equilibrium. The second is the maximum leeway, which instead computes the equilibrium under one-party control. The latter allows us to measure the strength of institutions separately from party control, which varies across states and over time. We emphasize that both of these leeway variables are summary measures of our high-dimensional treatment variables of interest (i.e., institutional features) and are not functions of the outcome variable of interest. Therefore, our methodology places an observational study of institutional changes under the design-based approach of causal inference (Rubin,, 2008).
@@
Specifically, we address the selection bias problem by using our leeway measures as a continuous treatment variable in a differences-in-differences (DiD) design applied to the 2010 and 2020 redistricting cycles. For example, states with Democratic control tend to enact more liberal policies (Caughey et al.,, 2017). This approach addresses potential confounding by comparing changes in states that have enacted reforms to those in similar states that have not, under a parallel trends assumption (Callaway et al.,, 2024). We use this model to estimate how changes in the map drawer’s leeway influence resulting plans across a set of both partisan and nonpartisan outcomes, such as the efficiency gap (Stephanopoulos and McGhee,, 2015) and responsiveness to swings in partisan preferences.
Our approach uses a formal model to summarize a multi-dimensional treatment variable, and then apply a causal inference methodology to adjust for bias due to unobserved confounding. However, we do not adopt a structural modeling approach, which would require modeling a state’s decision to adopt particular reforms for their redistricting process. The reason is that, while we are able to measure all the institutional features of redistricting processes, it is impossible to observe all the factors which affect a state’s decision to adopt redistricting reforms. Instead, we use a DiD design to address the issue of unobserved confounding.
@@
Lastly, we deal with the issue of outcome complexity by separating the causal effect of state-specific institutional reforms from changes in political geography over time. To do this, we generate a sample of alternative redistricting plans via a simulation algorithm (McCartan and Imai,, 2023) following each state’s rules and political geography, but without regard to any partisan information (McCartan et al.,, 2022; Kenny et al.,, 2024). We use these simulated plans as a nonpartisan baseline for both 2010 and 2020 redistricting cycles, and compute the difference in outcome variables between the enacted and simulated plans. This adjusts for state-specific changes in political geography, making the parallel trends assumption more plausible.
@@
We find that more restrictive redistricting processes reduce partisan bias by constraining map drawers. For example, changing from a single-party legislature to an independent commission leads to a reduction of about 0.5 excess seats. These effects are substantively large, as congressional gerrymandering generally results in gains of less than one or two seats per state (Kenny et al.,, 2023). Similarly, we find smaller but positive effects of constraining reforms on electoral responsiveness. We estimate that a similar change from legislature to commission would increase the share of competitive seats within a state from 25% to 38% on average.
@@
A key advantage of our methodological approach is the ability to perform a counterfactual analysis of institutional reforms. We investigate how enacting recent procedural reforms nationwide could reduce widespread partisan gerrymandering. We quantify how the partisan bias and responsiveness of adopted plans could counterfactually change if all states nationwide adopted three kinds of commission structures currently enacted in several states: (1) an Ohio-style approach that requires supermajorities and uses a bipartisan backup commission, (2) a New York-style commission with a nonpartisan map drawer but several partisan veto points later in the process, and (3) a Michigan-style commission with a nonpartisan commission drawer, no partisan veto points, and the potential for court review.
@@
We find that commissions can generally reduce existing pro-Republican bias, but the details of commission structure matter. In particular, unlike reforms adopted in Ohio and New York, a Michigan-style nonpartisan commission has no partisan veto points. We find that implementing Michigan-style reforms nationwide leads to an additional 22 Democratic seats, on average. Michigan-style reforms also increase electoral responsiveness whereas those adopted in New York and Ohio do not.
@@
Beyond the substantive findings, this paper advances a new methodology for studying redistricting reform, which we hope will translate to the study of other institutional systems. Reform efforts in redistricting have produced diverse institutional changes across states. While most scholars have classified reforms into different categories (Cain,, 2012; Edwards et al.,, 2017; Warshaw et al.,, 2022; Nelson,, 2023), such an approach may miss important differences in these institutional changes.
@@
In contrast, we use formal modeling to place these complex institutional characteristics on a continuous univariate scale to summarize how they constrain partisan actors. This theoretically driven approach, which we demonstrate accurately predicts empirical patterns, makes it possible to apply a difference-in-differences strategy to our complex setting for credible causal inference.
@@
While our work estimates the causal effects of redistricting reform as a whole, much of existing causal analysis has focused on one part of the process, such as voter registration patterns after the Shelby decision (Komisarchik and White,, 2021) and turnout in packed districts (Fraga et al.,, 2022). Other studies have reported positive associations between the adoption of commissions and competitiveness (Carson and Crespin,, 2004; Carson et al.,, 2014; Best et al.,, 2021; Nelson,, 2023) as well as a negative association between commissions and incumbent protection (Nelson,, 2023). These studies, however, do not adjust for various confounding factors considered in our analysis.
@@
Beyond redistricting, our methodological approach can be seen as a general strategy for the causal analysis of complex institutional reforms. By combining a game-theoretic treatment model with causal inference methods, we are able to leverage the strengths of these two approaches. A game-theoretic approach is a substantively effective way to model formal institutions with specific rules, making it possible to map multidimensional policies to a univariate summary in a theoretically-informed way. Once this summarization of various institutional reforms is done, we can apply standard causal inference methods to estimate the effects of counterfactual policies by mapping them directly to this univariate treatment variable.
@@
Our work relates to a broader methodological literature in the social sciences. In economics, for example, Chetty, (2009) advocates the combined use of structural and reduced-form approaches via a sufficient statistic in a manner similar to our approach. In sociology, Lundberg et al., (2021) emphasize the importance of theoretically informed quantities of interest in causal analysis. Finally, in political science, Canen and Ramsay, (2023) call for the integration of rigorous theoretical and empirical approaches in causal research. We demonstrate how such an analysis can be done in the estimation of causal effects of redistricting reforms.
@@ 2 Overview of Redistricting Processes and Reforms
Before introducing our methodological framework, we provide a brief overview of redistricting processes and reforms. Every decade following the U.S. Census, states and localities use many different procedures to redraw their legislative district boundaries. For example, states differ on whether legislatures or independent redistricting commissions propose initial plans for new Congressional district maps. If these actors fail to produce a plan, state laws further vary on how these stalemates are handled. Some states pass responsibilities to a court (e.g., Virginia), a backup commission (e.g., Ohio), or a group of state party leaders (e.g., Iowa). Even plans that pass the proposal step can face a veto from other actors, such as the governor and state legislature.
As explained in Section 3.1 below, we collect information about the redistricting process used in each state for the 2010 and 2020 cycles. Figure 1 summarizes this dataset and illustrates the diversity of 2020 redistricting procedures across states. For example, 29 states drew initial plans in their legislature, while 15 states used an independent redistricting commission instead. There is significant variation in the procedures following this initial draw, with seven distinct veto mechanisms across all the states. Much of this variation originates from state-specific policy campaigns that have adopted widely different goals driven by local actors (Keena et al.,, 2021). For example, California created a redistricting commission that has the power of drawing congressional district boundaries through a series of ballot propositions in the early 2000s. In recent years, several states have also implemented redistricting reforms in response to organized efforts to limit partisan influence. For example, Michigan voters approved a constitutional amendment to establish an independent redistricting commission in 2018.
@@
While these simplifications make the study of complex procedures tractable, they risk oversimplifying complex procedures in two ways. First, attributing institutional outcomes to a single actor ignores the fact that the final redistricting plan is produced through a series of steps. Take, for example, the New York redistricting process in 2020. An independent bipartisan commission had the power to propose an initial plan, but failed to agree on a single plan. State law required the stalemated process to move to the legislature, which adopted a plan that Republican and civil rights groups criticized as a Democratic gerrymander. These groups challenged the adopted plan in a series of lawsuits, and in 2022 the New York Court of Appeals struck down the plan and tasked a court-appointed special master with drawing a remedial plan. Though the final plan scores well on quantitative fairness metrics (Kenny et al.,, 2023), simply classifying a “court” as the sole map drawer for New York in 2020 ignores the partisan interests involved, failing to capture the complexity of the procedures that led to the final plan.
Second, classification schemes can overlook strategic interactions, where the behavior of certain actors depends on the presence or characteristics of others. For example, a commission might draw a different map if it knew that the map could later be reviewed by a court. Or, the potential of a governor’s veto may limit the likelihood of a partisan gerrymander by the state legislature, but not in cases where the governor and legislative majority share a partisan affiliation. More detailed coding of procedural schemes can account for some of these interaction effects and increase realism, but will necessarily decrease statistical power, making it more difficult to estimate causal effects.
Thus, we face a methodological dilemma: while some simplification of institutional features is necessary, common approaches to doing so obscure critical characteristics of redistricting processes. In the next section, we propose a theoretically grounded approach that models redistricting as a zero-sum sequential game and uses its Nash equilibrium as a treatment variable to summarize these complex institutions.
@@ 3 A Theoretical Model of Institutional Leeway
To study the impact of redistricting reforms and processes, we first develop a standardized set of 14 institutional features that capture the most important actors in congressional redistricting processes across states. We collect data on these features so that it is possible to compare redistricting processes consistently across states, despite the differences in the laws and bodies governing redistricting in each state.
Next, we develop a sequential redistricting game to summarize these features in a theoretically informed manner. Specifically, we use the Nash equilibrium of this game as a one-dimensional measure of the institutional leeway political actors have over the partisan lean of the resulting redistricting plan. Finally, we empirically validate the proposed measure by demonstrating that it is not particularly sensitive to model specification and predicts redistricting outcomes well. We emphasize that the proposed measure is a theoretically motivated summary of our high-dimensional treatment variable of interest. Since this measure does not use the outcome variable, we are able to apply the standard causal inference methodology to adjust for observed and unobserved confounding factors.